# Surface Reconstruction from Point Clouds via Grid-based Intersection Prediction

### 21 Mar 2024 v1

@@ 3.5 Intersection Module
Next, we can delve into the explanation for why a surface may not exhibit continuity if $H(f_{A},f_{B})\neq 1-H(f_{B},f_{A})$. Referring to Fig. 2 (a), the black cubes represent adjacent cubes in the Marching Cube algorithm. It is evident that $A$ is equivalent to $C$ and $B$ is equivalent to $D$. If a mesh is depicted by the red line, points $P$ and $Q$ must coincide; otherwise, the mesh will lack continuity. Therefore, the relationship $\frac{|AP|}{|AB|}=1-\frac{|DQ|}{|CD|}$ must be upheld. This condition serves as both necessary and sufficient for ensuring $H(f_{A},f_{B})=1-H(f_{B},f_{A})$.
3.6 Losses
@@
$Loss_{1}=\sum_{j=1}^{M}[\alpha_{j}-\alpha_{j_{GT}}]_{+},$ (3)
@@
$[\alpha-\alpha_{GT}]_{+}=\left\{\begin{array}[]{lc}|\alpha-\alpha_{GT}|&\alpha_{GT}\in[0,1]\\ 0&\alpha_{GT}\notin[0,1].\\ \end{array}\right.$ (4)
4 Experiments
4.1 Dataset and Experiment Settings
@@
Figure 6: Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR[11], CAP[31], GIFS[28], SuperUDF[10]. The number of input point cloud is 3000.
@@
Our methodology is assessed on three prominent open-source datasets. Initially, we randomly select 1300 shapes from 13 distinct categories within the ShapeNet dataset [5], with each category comprising 100 shapes. The point number of each input point cloud is 3000. Subsequently, we employ the MGN dataset, encompassing both pants and shirts. The point number of each input point cloud is 3000. Finally, we include 100 scenes from the real-world ScanNet dataset [8], obtained through scanning technology, resulting in 10,000 input point clouds. To facilitate a comprehensive comparison with existing techniques, we utilize standard evaluation metrics commonly employed in surface reconstruction. One such metric is Chamfer-$L_{1}$ ($CD_{1}$), which measures the distance between two sets of points. Additionally, to demonstrate enhancements in artifact elimination, we utilize normal consistency ($NC$) as a metric to evaluate the smoothness of the reconstructed surface.
@@
When calculating $CD_{1}$, we follow O-Net[19] and randomly sample 100000 points on the reconstructed mesh and 100000 points on ground-truth mesh. $CD_{1}$ equation is shown as the Eq.5
$\mathrm{CD}_{1}=$ $\frac{1}{2N_{x}}\sum_{i=1}^{N_{x}}\left\|\mathbf{x}_{i}-\mathcal{S}_{y}\left(\mathbf{x}_{i}\right)\right\|_{1}+$ (5)
@@
$\mathrm{NC}=$ $\frac{1}{2N_{x}}\sum_{i=1}^{N_{x}}|<\mathbf{n}\left(\mathbf{x}_{i}),\mathbf{n}(\mathcal{S}_{y}\left(\mathbf{x}_{i}\right)\right)>|+$ (6)
@@
4.2 Watertight Surface Reconstruction
@@
Table 1: $CD_{1}$ and $NC$ comparison on ShapeNet within 13 classes. $\downarrow$ means lower is better. $\uparrow$ means higher is better.
Several previous studies [21, 22, 19, 7] have commonly utilized the watertight surface reconstruction approach for implicit surface reconstruction, a key focus of this point cloud reconstruction research. In order to evaluate the effectiveness of our proposed method, we conducted a comprehensive analysis of watertight surface reconstruction using the ShapeNet dataset [5]. Specifically, we uniformly sampled 3000 points from the watertight mesh to serve as input data. We compared our method against four well-known techniques, excluding PSR [11], all of which are UDF-based methods. The methods included GIFS [28], CAP [31], and SuperUDF [10]. For the PSR [11] method, we incorporated ground-truth normals as input data. The visualization results are presented in Fig. 6. Our method not only preserves more details but also produces smoother results. The quantitative evaluation results are summarized in Table 1. These results demonstrate that our method achieves excellent performance across all 13 classes with a significant margin, showcasing its state-of-the-art capabilities. Particularly noteworthy is the performance on the $NC$, indicating that our reconstructed surfaces are not only faithful to the ground truth but also exhibit smoothness.
4.3 Open Surface Reconstruction
Figure 7: Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR[11], CAP[31], GIFS[28], SuperUDF[10]. The number of input point cloud is 3000.
CAP [31] PSR [11] GIFS [28] SUDF[10] Ours
@@ 4.3 Open Surface Reconstruction
Table 2: $CD_{1}$ and $NC$ comparison results on MGN dataset.
By leveraging a pair-point structure akin to GIFS for representing surfaces within our framework, we are able to proficiently reconstruct open surfaces. We achieve this by uniformly sampling 3000 points on each MGN mesh to reconstruct the implicit surface, a task that poses challenges for SDF-based methods given the characteristics of open surfaces. The comparative analysis presented in Table 2 illustrates the superior performance of our approach in comparison to various methods including PSR [11], CAP [31], GIFS [28], and SuperUDF [10]. Moreover, the visual results depicted in Fig.  7 exhibit the faithfulness of our reconstructed mesh, capturing intricate details and closely mirroring the ground-truth.
4.4 Scene Reconstruction
Real scene reconstruction is characterized by a heightened level of complexity, primarily due to the inherent openness and incompleteness of the input point cloud data. In this study, we meticulously selected 100 scenes from ScanNet and illustrated the visualization results in Fig. 8, along with the quantitative results in Table 3, in comparison with other methods.
@@ 4.4 Scene Reconstruction
CAP [31] PSR [11] GIFS [28] SuperUDF [10] Ours
@@
Table 3: $CD_{1}$ comparison on ScanNet with different number of sample point.
Figure 8: Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR[11], CAP[31], GIFS[28], SuperUDF[10]. The number of input point cloud is 3000.
4.5 Ablation Study
Contribution of sign module and intersection module.
To enhance the quality of surface reconstruction, it is essential to predict the sign of cube corners. This allows for the selection of appropriate triangle templates akin to the Marching Cube algorithm. Additionally, accurate intersection point prediction is crucial. By leveraging this information, adjustments can be made to the triangle positions, resulting in smoother surfaces with reduced artifacts. Our approach consists of two key components: the sign module and the intersection module. Consequently, we aim to conduct experiments to evaluate the impact of each component on surface quality. In pursuit of this objective, we have devised 9 comparative experiments to demonstrate the contributions of each component, as presented in Table 4 and Table 5. Upon examining the results row-wise, it is evident that regardless of the chosen sign prediction method, our intersection results outperform those of GIFS [28]. Furthermore, the disparity between our intersection prediction outcomes and ground truth (GT) predictions is minimal, indicating the enhanced accuracy of our intersection module. When analyzing the results column-wise, it is apparent that regardless of the chosen intersection prediction method, our sign prediction results surpass those of GIFS [28]. Moreover, the difference between our sign prediction outcomes and GT sign predictions is negligible, highlighting the efficacy of our sign module in improving sign prediction accuracy.
@@ 4.5 Ablation Study
Table 4: The contribution of sign module and intersection module. The $CD_{1}$ is obtained on ShapeNet 13 classes.
@@
Table 5: The contribution of sign module and intersection module. The $NC$ is obtained on ShapeNet 13 classes.
@@
Method GIFS [28] SuperUDF[10] CAP [31] Ours
@@
Table 6: The distance from predicted intersection to GT intersection on ShapeNet.
@@
Table 7: The reconstruction result comparison when sign module is invariant or variant to the exchange of start and end point. The result is average on 13 classes of ShapeNet.
@@
Table 8: The Comparison when intersection module satisfies $H(f_{A},f_{B})=1-H(f_{B},f_{A})$ or not. The result is average on 13 classes of ShapeNet.
Sign module in-variance.
In our design, the sign module remains invariant when the start point and end point exchange their positions, i.e., $G(f_{A},f_{B})=G(f_{B},f_{A})$. Here, we aim to quantitatively evaluate the influence of this design feature. Therefore, we compare the reconstruction results when $G(\cdot,\cdot)$ does not maintain this invariance with the results when $G(\cdot,\cdot)$ does maintain this invariance. Specifically, when $G(\cdot,\cdot)$ does not maintain the invariance, we set
$\hat{G}(f_{A},f_{B})=concat(f_{A}+pos_{3},f_{B}+pos_{3}),$ (7)
where $pos_{3}$ is the standard positional encoding of both sine and cosine encoding. The quantitative result is in Table 7. We can see the invariance works.
@@
Figure 9: When the intersection module satisfy $H(f_{A},f_{B})=1-H(f_{B},f_{A})$, the surface is continuous, otherwise it is not continuous and has lots of errors.
5 Conclusion and Limitations
This study introduces a method that predicts the intersection between a segment of a point pair and an implicit surface to effectively eliminate artifacts. Additionally, it proposes two novel modules that utilize neural networks to predict the sign of cube corners and intersections. This approach results in a more detailed surface reconstruction with fewer artifacts, leading to excellent outcomes on three datasets: ShapeNet, MGN, and ScanNet. However, there are some limitations to this method. It performs accurately when the surface is parallel to the bounding box face; yet, significant changes in the angle between the tangent plane of the surface and the bounding box face in a local region can lead to errors in intersection prediction, resulting in a less smooth surface. This issue will be addressed in future research. Furthermore, the method struggles to generate satisfactory meshes near the edges of open surfaces.
Appendix 0.A Details
Training.
During training, we utilize 16,000 randomly sampled point pairs every iteration. The trainig epoch is 100. We employ the Adam optimizer [12] and incorporate cosine learning rate adjustment [16]. The initial learning rate is set to 0.001, and the batch size is 1.
Testing.
Training/Testing splitting.
Two prediction ways.
Similarities and differences between GIFS [28].
Post-processing.
$I=\frac{1}{k}\sum_{i=0}^{k-1}|<n,n_{i}>|,$ (9)
Appendix 0.B Additional visualization result
In order to show the quality of reconstructed mesh, we provide additional visualization result in Fig. 10, Fig. 11 and Fig. 12.
Figure 10: Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR[11], CAP[31], GIFS[28], SuperUDF[10]. The number of input point cloud is 3000.
Figure 11: Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR[11], CAP[31], GIFS[28], SuperUDF[10]. The number of input point cloud is 3000.
Figure 12: Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR[11], CAP[31], GIFS[28], SuperUDF[10]. The number of input point cloud is 10000.
Appendix 0.C Efficiency Analysis
@@
Table 9: The time and memory comparison between GIFS [28], CAP [31] and ours in inference. The result is on the ShapeNet.

### 09 Apr 2024 v2

@@ 3.5. Intersection Module
Next, we can delve into the explanation for why a surface may not exhibit continuity if $H(f_{A},f_{B})\neq 1-H(f_{B},f_{A})$. Referring to Fig. 6, the two cubes represent adjacent cubes in the Marching Cube algorithm. It is evident that $A$ is equivalent to $C$ and $B$ is equivalent to $D$. Pints $P$ and $Q$ must coincide; otherwise, the mesh will lack continuity. Therefore, the relationship $\frac{|AP|}{|AB|}=1-\frac{|DQ|}{|CD|}$ must be upheld. This condition serves as both necessary and sufficient for ensuring $H(f_{A},f_{B})=1-H(f_{B},f_{A})$ and $G(f_{A},f_{B})=G(f_{B},f_{A})$.
3.6. Losses
@@
(3) $Loss_{1}=\sum_{j=1}^{M}[\alpha_{j}-\alpha_{j_{GT}}]_{+},$
@@
(4) $[\alpha-\alpha_{GT}]_{+}=\left\{\begin{array}[]{c}|\alpha-\alpha_{GT}|,\alpha_{GT}\in[0,1]\\ 0,\alpha_{GT}\notin[0,1].\\ \end{array}\right.$
3.7. Details and Discussions
Training.
During training, we utilize 16,000 randomly sampled point pairs every iteration. The training epoch is 100. We employ the Adam optimizer (Kingma et al., 2014) and incorporate cosine learning rate adjustment (Loshchilov and Hutter, 2016). The initial learning rate is set to 0.001, and the batch size is 1.
Testing.
Training/Testing splitting.
Two prediction ways.
Figure 6. The symmetry ensures the continuity of surface.
Post-processing.
(5) $I=\frac{1}{k}\sum_{i=0}^{k-1}|<n,n_{i}>|,$
Similarities and differences between GIFS (Ye et al., 2022) and Ours.
Similarities and differences of Deep Marching Cube, Neural Marching Cube and Ours.
Deep Marching Cube (Liao et al., 2018) is a work closely related to ours, as both focus on the two key aspects: the sign of vertices at the corner of the cube and the position of intersection points. However, several differences distinguish our approaches. Firstly, our motivation differs. While our goal is to eliminate artefacts, Deep Marching Cube aims to differentiate the Marching Cube. Secondly, our method is UDF-based, allowing us to represent open surfaces, whereas Deep Marching Cube, based on SDF, can only represent watertight surfaces. Thirdly, we directly predict the relative sign of point pairs, while Deep Marching Cube employs a probabilistic model to represent vertex signs.
Neural Marching Cube (Chen and Zhang, 2021) seeks to enhance Marching Cube performance by intersecting additional vertices beyond those on the cube edges, aiming to generate a finer mesh without cube subdivision. Consequently, our approach and Neural Marching Cube employ different methods to address distinct problems, indicating a somewhat distant relationship between them.
4. Experiments
4.1. Tiny example to show the difficulty of learning UDF
Compared to SDF, UDF does not require predicting signs, making it seem simpler than SDF. However, this is not the case. We designed a small experiment to demonstrate the difficulties of learning UDF, using a simple neural network to fit the circle function on a two-dimensional plane. Our goal is to learn the UDF and SDF of a circle on the plane, using a three-layer MLP (Multi-Layer Perception) to represent this sphere, with optimization done through gradient descent. We provide two three-layer MLP functions $\mathcal{F}_{\theta_{1}}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{+}$ to represent the UDF field on the plane, and $\mathcal{F}_{\theta_{2}}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ to represent the SDF field on the plane. The parameters of the experiment are as follows: the output dimensions of the three-layer MLP are $32,64,1$ respectively, the number of sampled points is $20000$, the initial learning rate is $0.001$, and the optimizer used is Adam. The visualization of the learning results is shown in Fig. 7.
From the results of SDF and UDF in the first row, it can be seen that the neural network indeed learns the approximate distance field. When the distance field changes along the x-axis at the position of $y=0$ in the second row, we focus on the error between the GT distance field and the learned distance field to the circle as x changes. It can be observed that when learning SDF, the error between the two is almost zero, while when learning UDF, the error is significant. This is because near the surface of the curve, the gradient of the SDF remains unchanged, resulting in a continuous change in the distance field, while the gradient of the UDF rotates by $180^{\circ}$, making the change in the distance field discontinuous. Specifically, when x approaches $-1$ from the left side, the gradient of the SDF is
(6) $\lim_{x\rightarrow 0^{+}}\nabla\mathcal{F}_{\theta_{2}}=(-1,0)$
When x approaches -1 from the right, the gradient of the SDF is
(7) $\lim_{x\rightarrow 0^{-}}\nabla\mathcal{F}_{\theta_{2}}=(-1,0)$
They are same with each other in SDF. when x approaches $-1$ from the left side, the gradient of the UDF is
(8) $\lim_{x\rightarrow 0^{+}}\nabla\mathcal{F}_{\theta_{1}}=(-1,0)$
When x approaches -1 from the right, the gradient of the UDF is
(9) $\lim_{x\rightarrow 0^{-}}\nabla\mathcal{F}_{\theta_{1}}=(1,0)$
The two are exactly opposite. Although the Signed Distance Function (SDF) also has some errors at $x=0$ due to the collision of gradient directions, it is insignificant because it occurs at the center of the sphere, which is not important for surface reconstruction. Therefore, the error here can be ignored. However, the Unsigned Distance Function (UDF) has a discontinuity in the vicinity of the surface, and the distance field near the surface is particularly important for surface reconstruction. For regression methods, fitting a continuously changing function is relatively easy, while fitting a function with gradient discontinuities is very difficult. This is the difficulty of reconstruction based on UDF. The difficulty of reconstruction based on SDF is easier to understand, mainly in complex topological surfaces where it is difficult for neural networks to determine the sign information at a specific point.
Figure 7. Comparison of learning difficulties between UDF and SDF. Using neural networks to learn the Signed Distance Function (SDF) and Unsigned Distance Function (UDF) from any point in the plane to the circle represented by the green line in the first row of the image. The second row shows the variation of SDF and UDF along the x-axis from the position at $y=0$.
4.2. Dataset and Experiment Settings
@@
Figure 8. Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR(Kazhdan et al., 2013), CAP(Zhou et al., 2022), GIFS(Ye et al., 2022), SuperUDF(Tian et al., 2023). The point number of input point cloud is 3000.
@@
Our methodology is assessed on three prominent open-source datasets. Initially, we randomly select 1300 shapes from 13 distinct categories within the ShapeNet dataset (Chang et al., 2015), with each category comprising 100 shapes. The point number of each input point cloud is 3000. Subsequently, we employ the MGN dataset, encompassing both pants and shirts. The point number of each input point cloud is 3000. Finally, we include 100 scenes from the real-world ScanNet dataset (Dai et al., 2017), obtained through scanning technology, resulting in 10,000 input point clouds. To facilitate a comprehensive comparison with existing techniques, we utilize standard evaluation metrics commonly employed in surface reconstruction. One such metric is Chamfer-$L_{1}$ ($CD_{1}$), which measures the distance between two sets of points. Additionally, to demonstrate enhancements in artifact elimination, we utilize normal consistency ($NC$) as a metric to evaluate the smoothness of the reconstructed surface.
@@
When calculating $CD_{1}$, we follow O-Net(Mescheder et al., 2019) and randomly sample 100000 points on the reconstructed mesh and 100000 points on ground-truth mesh. $CD_{1}$ equation is shown as the Eq.10
(10) $\mathrm{CD}_{1}=$ $\frac{1}{2N_{x}}\sum_{i=1}^{N_{x}}\left\|\mathbf{x}_{i}-\mathcal{S}_{y}\left(\mathbf{x}_{i}\right)\right\|_{1}+$
@@
(11) $\mathrm{NC}=$ $\frac{1}{2N_{x}}\sum_{i=1}^{N_{x}}|<\mathbf{n}\left(\mathbf{x}_{i}),\mathbf{n}(\mathcal{S}_{y}\left(\mathbf{x}_{i}\right)\right)>|+$
@@
4.3. Watertight Surface Reconstruction
@@
Table 1. $CD_{1}$ and $NC$ comparison on ShapeNet within 13 classes. $\downarrow$ means lower is better. $\uparrow$ means higher is better. Comparing methods are GIFS (Ye et al., 2022),CAP (Zhou et al., 2022),PSR (Kazhdan et al., 2013),SuperUDF (Tian et al., 2023). The point number of input point cloud is 3000.
Several previous studies (Park et al., 2019; Peng et al., 2020; Mescheder et al., 2019; Chibane et al., 2020b) have commonly utilized the watertight surface reconstruction approach for implicit surface reconstruction, a key focus of this point cloud reconstruction research. In order to evaluate the effectiveness of our proposed method, we conducted a comprehensive analysis of watertight surface reconstruction using the ShapeNet dataset (Chang et al., 2015). Specifically, we uniformly sampled 3000 points from the watertight mesh to serve as input data. We compared our method against four well-known techniques, excluding PSR (Kazhdan et al., 2013), all of which are UDF-based methods. The methods included GIFS (Ye et al., 2022), CAP (Zhou et al., 2022), and SuperUDF (Tian et al., 2023). For the PSR (Kazhdan et al., 2013) method, we incorporated ground-truth normals as input data. The visualization results are presented in Fig. 8 and Fig.  9. Our method not only preserves more details but also produces smoother results. One of the reasons for the unsatisfactory performance of CAP (Zhou et al., 2022) is the insufficient number of input points in the point cloud. With only 3000 points available for training, CAP typically requires over 20000 points in the point cloud to achieve optimal results. Typically, methods such as ConvOccNet(Peng et al., 2020) and POCO(Boulch and Marlet, 2022) initially learn features from sparse point clouds and then decode the distance field of query points. Conversely, approaches like Neural Pull(Ma et al., 2020) and CAP(Zhou et al., 2022) train neural networks to approximate the distance field, often necessitating dense point clouds. Our method falls within the former category.
The quantitative evaluation results are summarized in Table 1. These results demonstrate that our method achieves excellent performance across all 13 classes with a significant margin, showcasing its state-of-the-art capabilities. Particularly noteworthy is the performance on the $NC$, indicating that our reconstructed surfaces are not only faithful to the ground truth but also exhibit smoothness.
Figure 9. Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR(Kazhdan et al., 2013), CAP(Zhou et al., 2022), GIFS(Ye et al., 2022), SuperUDF(Tian et al., 2023). The point number of input point cloud is 3000.
4.4. Open Surface Reconstruction
Figure 10. Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR(Kazhdan et al., 2013), CAP(Zhou et al., 2022), GIFS(Ye et al., 2022), SuperUDF(Tian et al., 2023). The point number of input point cloud is 3000.
CAP PSR GIFS SuperUDF Ours
@@ 4.4. Open Surface Reconstruction
Table 2. $CD_{1}$ and $NC$ comparison results on MGN dataset. Comparing methods are GIFS (Ye et al., 2022),CAP (Zhou et al., 2022),PSR (Kazhdan et al., 2013),SuperUDF (Tian et al., 2023). The point number of input point cloud is 3000.
By leveraging a pair-point structure akin to GIFS for representing surfaces within our framework, we are able to proficiently reconstruct open surfaces. We achieve this by uniformly sampling 3000 points on each MGN mesh to reconstruct the implicit surface, a task that poses challenges for SDF-based methods given the characteristics of open surfaces. The comparative analysis presented in Table 2 illustrates the superior performance of our approach in comparison to various methods including PSR (Kazhdan et al., 2013), CAP (Zhou et al., 2022), GIFS (Ye et al., 2022), and SuperUDF (Tian et al., 2023). Moreover, the visual results depicted in Fig.  10 and Fig.  11. exhibit the faithfulness of our reconstructed mesh, capturing intricate details and closely mirroring the ground-truth.
Figure 11. Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR(Kazhdan et al., 2013), CAP(Zhou et al., 2022), GIFS(Ye et al., 2022), SuperUDF(Tian et al., 2023). The point number of input point cloud is 3000.
4.5. Scene Reconstruction
Real scene reconstruction is characterized by a heightened level of complexity, primarily due to the inherent openness and incompleteness of the input point cloud data. In this study, we meticulously selected 100 scenes from ScanNet and illustrated the visualization results in Fig. 13 and Fig. 12, along with the quantitative results in Table 3, in comparison with other methods.
@@ 4.5. Scene Reconstruction
Figure 12. Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR(Kazhdan et al., 2013), CAP(Zhou et al., 2022), GIFS(Ye et al., 2022), SuperUDF(Tian et al., 2023). The point number of input point cloud is 10000.
CAP PSR GIFS SuperUDF Ours
@@
Table 3. $CD_{1}$ and $NC$ comparison on ScanNet. Comparing methods are GIFS (Ye et al., 2022),CAP (Zhou et al., 2022),PSR (Kazhdan et al., 2013),SuperUDF (Tian et al., 2023). The point number of input point cloud is 10000.
Figure 13. Visualized results of our method and state-of-the-art alternatives. Note that we provide the ground-truth normal as input for PSR as it requires. Methods are PSR(Kazhdan et al., 2013), CAP(Zhou et al., 2022), GIFS(Ye et al., 2022), SuperUDF(Tian et al., 2023). The point number of input point cloud is 10000.
4.6. Ablation Study
Contribution of Relative Sign Module and Intersection Module.
To enhance the quality of surface reconstruction, it is essential to predict the sign of cube corners. This allows for the selection of appropriate triangle templates akin to the Marching Cube algorithm. Additionally, accurate intersection point prediction is crucial. By leveraging this information, adjustments can be made to the triangle positions, resulting in smoother surfaces with reduced artefacts. Our approach consists of two key components: the Relative Sign Module and the Intersection Module. Consequently, we aim to conduct experiments to evaluate the impact of each component on surface quality. In pursuit of this objective, we have devised 9 comparative experiments to demonstrate the contributions of each component, as presented in Table 4 and Table 5. Upon examining the results row-wise, it is evident that regardless of the chosen sign prediction method, our intersection results outperform those of GIFS (Ye et al., 2022). Furthermore, the disparity between our intersection prediction outcomes and ground truth (GT) predictions is minimal, indicating the enhanced accuracy of our Intersection Module. When analyzing the results column-wise, it is apparent that regardless of the chosen intersection prediction method, our sign prediction results surpass those of GIFS (Ye et al., 2022). Moreover, the difference between our sign prediction outcomes and GT sign predictions is negligible, highlighting the efficacy of our Relative Sign Module in improving sign prediction accuracy.
@@ 4.6. Ablation Study
Table 4. The contribution of Relative Sign Module and Intersection Module. The $CD_{1}$ is obtained on ShapeNet 13 classes.
@@
Table 5. The contribution of Relative Sign Module and Intersection Module. The $NC$ is obtained on ShapeNet 13 classes.
@@
Method GIFS SuperUDF CAP Ours
@@
Table 6. The distance from predicted intersection to GT intersection on ShapeNet. Comparing methods are GIFS (Ye et al., 2022),CAP (Zhou et al., 2022), SuperUDF (Tian et al., 2023).
@@
Table 7. The reconstruction result comparison when Relative Sign Module is symmetrical or asymmetrical to the exchange of start and end point. The result is averaged on 13 classes of ShapeNet.
@@
Table 8. The Comparison when Intersection Module satisfies $H(f_{A},f_{B})=1-H(f_{B},f_{A})$ or not. The result is average on 13 classes of ShapeNet.
Relative Sign Module Symmetry.
In our design, the Relative Sign Module remains symmetrical when the start point and end point exchange their positions, i.e., $G(f_{A},f_{B})=G(f_{B},f_{A})$. Here, we aim to quantitatively evaluate the influence of this design feature. Therefore, we compare the reconstruction results when $G(\cdot,\cdot)$ does not maintain this symmetry with the results when $G(\cdot,\cdot)$ does maintain this symmetry. Specifically, when $G(\cdot,\cdot)$ does not maintain the symmetry, we set
(12) $\hat{G}(f_{A},f_{B})=concat(f_{A}+pos_{3},f_{B}+pos_{3}),$
where $pos_{3}$ is the standard positional encoding of both sine and cosine encoding. The quantitative result is in Table 7. We can see the symmetry works.
@@
Figure 14. When the Intersection Module satisfies symmetry, the surface is continuous, otherwise it is not continuous and has lots of errors.
Efficiency Analysis
@@
Table 9. The time and memory comparison between GIFS (Ye et al., 2022), CAP (Zhou et al., 2022) and ours in inference. The result is on the ShapeNet.
5. Conclusion and Limitations
This study introduces a method that predicts the intersection between a line segment of a point pair and an implicit surface to effectively eliminate artefacts. Additionally, it proposes two novel modules that utilize neural networks to predict the relative sign of vertices at the corners of the cube and intersections between surface and line segment. This approach results in a more detailed surface reconstruction with fewer artefacts, leading to excellent outcomes on three datasets: ShapeNet, MGN, and ScanNet. However, there are some limitations to this method. For example, the method struggles to generate satisfactory meshes near the edges of open surfaces. This is problem will be addressed in the future works.